Graph the containing point p and having slope m. p=(8,-6); m=0
thanks for the help!

Answer by Edwin Parker(36)

Two ways to do it.

First way:  Using point slope formula:

y - y1 = m(x - x1) where (x1,y1) = (8,-6)

y - (-6) = 0(x - 8)

   y + 6 = 0

       y = -6

Second way:  Since the slope is 0, the line is horizontal, so the
graph is the black line below:



and since it goes through the y-axis where y is -6, its equation 
is y = -6, because all points om that line have their y-coordinate 
as -6. So you can just ignore the x-coordinate.  

Edwin